# MIBench: A Comprehensive Benchmark for Model Inversion Attack and Defense

### v2

@@ 3.3 Toolbox
We implement an extensible and reproducible modular-based toolbox for our benchmark, as shown in Fig 1. The framework can be divided into four main modules, including Data Preprocess Module, Attack Module, Defense Module and Evaluation Module.
Data Preprocess Module. This module is designed to preprocess all data resources required before launching attacks or defenses, including datasets, classifiers and parameters. Consequently, we furnish this module with three fundamental functionalities: dataset preprocessing, target model training, and parameter management. For dataset preprocessing, we build a unified pipeline for each dataset, which automatically carries out a series of operations such as spliting dataset and image transformations (e.g. center crop and resize) based on the split file and chosen resolution from users. For target model training, users can further leverage the processed datasets to train designated classifiers. We abstract the general procedures in classifier training into a base trainer class to facilitate users in extending and customizing their own classifiers. For parameter management, we encapsulate parameters used in different processes into specific configuration classes, such as TrainConfig designated for the training process and AttackConfig for the attack process, thus ensuring organized and efficient parameter handling across various operations in the workflow.
Attack Module. The workflow of MI attacks can be roughly divided into two stages. The first stage is GAN training, where the module abstracts the general training process of GANs into a basic trainer class. This allows users not only train GANs that are pre-built into the benchmark, but also extend to their uniquely designed GANs. The second stage is the core inversion attack, which we split into three parts: latent vectors initialization, iterative optimization, and an optional post-processing step. After completion of the attack, the module preserves essential data such as the final optimized images and latent vectors, facilitating subsequent evaluation and analysis.
@@
Evaluation Module. Our benchmark concentrates on the evaluation at distribution level instead of sample level, assessing the overall performance of the whole reconstructed dataset. Therefore, we provide a total of 9 widely recognized distribution level evaluation metrics for users, which can be categorized into four types according to the evaluated content: Accuracy, Feature Distance, FID (Heusel et al., 2017) and Sample Diversity. For Accuracy, this metric measures how well the reconstructed samples resemble the target class, consisting of and . For Feature Distance, it computes the average shortest distance from features between each reconstructed sample and private data to measure the similarity between the feature space of reconstructed samples and private dataset. For FID, the lower FID score shows higher realism and overall diversity (Wang et al., 2021a). For Sample Diversity, the higher values indicate greater intra-class diversity. Moreover, we provide convenient tools for analysis, including Standard Deviation Calculation and Visualization.
4 Experiment
4.1 Experimental Setups
To ensure fair and uniform comparison and evaluation, we select the FFHQ (Karras et al., 2019) as the public dataset and FaceScrub (Ng & Winkler, 2014) as the private dataset for all the experiments in the Experiment section. The target models are fixed to the IR-152 (He et al., 2016) for low-resolution scenario and ResNet-152 (He et al., 2016) for high-resolution scenario, both trained on the FaceScrub. For each attack method, the number of images reconstructed per class is set to 5 due to considerations of time and computation cost. More detailed experimental settings are listed in Sec.B in the Appendix.
Notably, we limit the evaluation exhibited in the Experiment section to merely three metrics, including Accuracy, Feature Distance and FID (Heusel et al., 2017), while Sample Diversity is presented in the Sec.C.1 in the Appendix. Moreover, the VMI (Wang et al., 2021a) and RLBMI (Han et al., 2023) will be further evaluated in Sec. C.7 owing to their excessive need of time.
@@ 4.2 Evaluation on Different Attack Methods
In this part, we prepare a unified experimental setting for different MI attack methods to conduct a fair comparison. The resolution of private and public datasets is set to , indicating a relatively easier scenario. Comparisons of white-box and black-box MI attacks are presented in Table 3.
Remarkably, the PLGMI (Yuan et al., 2023a) and LOKT (Nguyen et al., 2023a) achieve state-of-the-art comprehensive performance in white-box attacks and black-box attacks respectively, showing significant superiority in Accuracy and Feature Distance metrics. However, the lowest FID scores occur in the PPA (Struppek et al., 2022) and C2FMI (Ye et al., 2023) respectively instead of the above methods. We infer that this is because PPA and C2FMI employ more powerful generators (e.g. StyleGAN2 (Karras et al., 2020b)) as the GAN prior compared to PLGMI and LOKT, leading to more real image generation. Visualization in Fig 2 further validates the inference.
@@ 4.3 Evaluation on Higher Resolution
The results imply the significant impact of GAN priors when attacking private images with higher resolution. All the methods that employ stronger GAN priors maintain low FID scores, including Mirror, C2FMI, and PPA, while other methods suffer from significant degradation in image reality. This phenomenon is more pronounced in the visualization results displayed in Fig 3. Despite the primary metric for evaluating MI attacks is Accuracy, the reality of reconstructed images should be ensured within a reasonable range for better image quality. Thus, it is imperative to explore more complex GAN priors with enhanced performance in future research, extending the MI field to more challenging and practical applications.
@@
4.4 Evaluation on Different Model Predictive Power
The predictive power of the target model is a crucial factor in determining the effectiveness of MI attacks. Previous work (Zhang et al., 2020a) has conducted preliminary experimental validations on simple networks (e.g. LeNet (LeCun et al., 1989)), demonstrating that the performance of the first GAN-based attack GMI is influenced by the predictive power of the target model. Therefore, we evaluate the state-of-the-art MI attacks on target models with varied predictive power to further validate the consistency of this characteristic in the recent attacks. The resolution of datasets is set to . The evaluation is stated in Fig. 4.
The comparison in Fig. 4 reveals most MI attacks maintain the trend that higher predictive power contributes to better attack performance, which is consistent with the aforementioned characteristic. Specifically, the earlier attack methods (e.g., GMI and KEDMI) presents more fluctuation on the trend across different predictive power, while the recent attack methods (e.g., PLGMI and PPA) show in more stable trend when predictive power increases. This indicates the predictive power of target models plays a crucial role in measuring the performance of MI attacks. Thus one can expect lower privacy leakage in the robust model by balancing the accuracy-privacy trade-off. This study of the predictive power of target models illustrates another useful aspect of our MIBench.
@@ 4.5 Evaluation on Defense Strategies
This analysis concentrates on the robustness of MI attacks when applied to target models with defense strategies. Notably, we select the first 100 classes subset from FaceScrub as the target dataset due to the time cost. The assessment results are listed in Fig.5. The configuration of defenses is set following the official parameters, as detailed in Table 8.
Overall, the TL (Ho et al., 2024) achieves the state-of-the-art decrease in Accuracy. However, advanced attack methods have overcome current defense strategies to some extent, such as PLG and PPA. Additionally, some older defense strategies (e.g. MID (Wang et al., 2021c)) are no longer effective against the latest attacks. From Fig.5, we observe that LS (Struppek et al., 2023) exhibits unexpectedly poor performance in metric while it was recently published in the top-tier conference. The potential reason might be the utilized target classifier with relatively high test accuracy, as validated in the Sec. C.4. Furthermore, we conduct further experiments with PPA as the attack method against ResNet-152 trained on high resolution (224×224) scenario, proving that this phenomenon can be extended into other defenses. The results are listed in the Table 5, demonstrating that all the defenses become invalid even with the recommended parameters. More in-depth evaluation on defenses is exhibited in Sec C.5.
Combining this phenomenon with the above experiment on predictive power, our empirical analysis indicates that the leaked information is strongly correlated to the model prediction accuracy and current defenses cannot effectively reduce the privacy information without sacrifice of model performance. Our findings emphasize that more reliable and stable defense strategies should be studied due to the fact that high model prediction accuracy is crucial for application of AI technology.
@@
| NO Defense | - | 98.510 | 0.972 | 0.990 | 307.714 | 0.588 | 50.259 |
| MID | 96.760 | 1.000 | 1.000 | 273.687 | 0.517 | 49.239 | |
| 95.680 | 0.990 | 1.000 | 276.889 | 0.526 | 51.227 | ||
| BiDO | 98.030 | 0.986 | 0.996 | 306.827 | 0.554 | 50.943 | |
| 97.430 | 0.968 | 0.996 | 320.191 | 0.594 | 50.132 | ||
| TL | 97.620 | 0.994 | 1.000 | 282.394 | 0.513 | 53.023 | |
| 97.530 | 0.982 | 0.996 | 306.528 | 0.561 | 51.489 |
@@ 5 Conclusion
In this paper, we develop MIBench, a comprehensive, unified, and reliable benchmark, and provide an extensible and reproducible toolbox for researchers. To the best of our knowledge, this is the first benchmark and first open-source toolbox in the MI field. Our benchmark encompasses 16 of the state-of-the-art MI attack methods and defense strategies and more algorithms will be continually updated. Based on the implemented toolbox, we establish a consistent experimental environment and conducted extensive experimental analyses to facilitate fair comparison between different methods. In our experiments, we explore the impact of multiple settings, such as different image resolutions, model predictive power and defense. With in-depth analysis, we have identified new insights and proposed potential solutions to alleviate them.
Societal Impact and Ethical Considerations. A potential negative impact of our benchmark could be malicious users leveraging the implemented attack methods to reconstruct private data from public system. To alleviate this potential dilemma, a cautious approach for data users is to adopt robust and reliable defense strategies, as shown in the Sec.4.5 of our paper. Additionally, establishing access permissions and limiting the number of visits for each user is crucial to build responsible AI systems, thereby alleviating the potential contradictions with individual data subjects.
@@ A.1 Details of Data Processing
This section introduces three types of dataset processing implemented by our toolbox, including data pre-processing, dataset splitting and dataset synthesis.
@@
FaceScrub (Ng & Winkler, 2014). In the low-resolution version, the original data is center cropped at and then scaled, and the high-resolution is scaled directly.
@@
This step is used to model training, including classifiers and conditional generators. For labeled datasets like CelebA and FaceScrub, we provide a fixed partition to slice the training and test sets. For unlabeled datasets such as FFHQ and MetFaces, we provide a script for pseudo-labeling all images by computing the scores. The strategy of PLGMI (Yuan et al., 2023a) for the score calculation is contained in the examples of our toolbox.
@@
Some methods use synthetic datasets for model training. We provide a script to synthesis datasets by pre-trained GANs according to LOKT (Nguyen et al., 2023a).
@@
We implemented and versions for all the custom attack models defined by the official code of implemented algorithms. For those algorithms that use StyleGAN2 (Karras et al., 2020b), we provide a wrapper for loading the official model and adapting it to our attack process.
@@ A.3 Implementation of Classifier Training & Defense Methods
MID (Wang et al., 2021c). It applies a Gaussian perturbation to the features in front of the classification head to reduce the mutual information of the model inputs and outputs. According to the official code, the method can also be called VIB, and we adopt this name in our toolbox. The hyperparameter controls the factor of the Gaussian perturbation term.
@@
BiDO (Peng et al., 2022). It adds the regular term loss (COCO or HSIC) function so that the intermediate features decrease the mutual information with the input and increase the mutual information with the output. It has two hyperparameters, and . The former is the factor of the regular term loss between inputs and features, and the latter means that between features and model outputs.
@@
Label Smoothing (LS) (Struppek et al., 2023). It adds an label smoothing term with a negative smoothing factor into the cross-entropy loss.
@@
Transfer Learning (TL) (Ho et al., 2024). This method uses a pre-trained model for fine-tuning. The parameters of the previous layers are frozen and those of the later layers will be fine-tuned. In our implementation, the hyperparameter is defined as a ratio of the number of frozen layers.

### v3

@@ 3.3 Toolbox
We implement an extensible and reproducible modular-based toolbox for our benchmark, as shown in Fig. 1. The framework can be roughly divided into four primary modules, including Data Preprocess Module, Attack Module, Defense Module and Evaluation Module.
Data Preprocess Module. This module is designed to preprocess all data resources required before launching attacks or defenses, including datasets, classifiers and parameters. Consequently, we furnish this module with three fundamental functionalities: dataset preprocessing, target model training, and parameter management. For dataset preprocessing, we build a unified pipeline for each dataset, which automatically carries out a series of operations such as splitting dataset and image transformations (e.g. center crop and resize) based on the split file and chosen resolution from users. For target model training, users can further leverage the processed datasets to train designated classifiers. We abstract the general procedures in classifier training into a base trainer class to facilitate users in extending and customizing their own classifiers. For parameter management, we encapsulate parameters used in different processes into specific configuration classes, such as TrainConfig, designated for the training process, and AttackConfig for the attack process, thus ensuring organized and efficient parameter handling in the workflow.
Attack Module. The workflow of MI attacks can be roughly divided into two stages. The first stage is GAN training, where the module abstracts the general training process of GANs into a basic trainer class. This allows users not only train GANs that are pre-built into the benchmark, but also extend to their uniquely designed GANs. The second stage is the core inversion attack, which we split into three parts: latent vectors initialization, iterative optimization, and an optional post-processing step. After the completion of the attack, the module preserves essential data such as the final optimized images and latent vectors, facilitating subsequent evaluation and analysis.
@@
Evaluation Module. Following the common setting in MI field, our benchmark concentrates on the evaluation at distribution level instead of sample level, assessing the overall performance of the whole reconstructed dataset. Therefore, we provide a total of 9 widely recognized distribution level evaluation metrics for users, which can be categorized into four types according to the evaluated content:
Attack Accuracy. This metric serves as the primary measurement for MI attacks, using a given classifier to classify the inverted images and measures the top- () and top- () accuracy for target labels. The higher the reconstructed samples achieve attack accuracy on another classifier trained with the same dataset, the more private information in the dataset can be considered to be exposed [63]. Therefore, higher attack accuracy denotes better attack performance.
Feature Distance. The feature is defined as the output of the given classifier’s penultimate layer. We compute the shortest feature distance between inverted images and private samples for each class and calculate the average distance. In our benchmark, we follow the common choices across previous methods [43, 39] and select the evaluation model and another pre-trained FaceNet [40] to compute the feature distance, respectively denoted as and . Smaller feature distance means more similar features to the private image.
Fréchet Inception Distance (FID). FID [19] is commonly used to evaluate the generation quality of GANs by computing the distance between feature vectors from all the target private data and that from all the reconstructed samples. Feature vectors are extracted by an Inception-v3 [46] pre-trained on ImageNet [7]. Lower FID score shows higher inter-class diversity and realism [49].
Sample Diversity. The metric contains Precision-Recall [27] and Density-Coverage [32] scores. Higher values indicate greater intra-class diversity of the inverted images. Since this metric serves as a secondary indicator for auxiliary reference in MI tasks, we have placed the corresponding evaluation results in the Supp.C.1.
4 Evaluation
4.1 Evaluation Setups
To ensure fair and uniform comparison and evaluation, we select the FFHQ [22] as the public dataset and FaceScrub [33] as the private dataset for all the experiments in the Experiment section. The target models are fixed to the IR-152 [17] for low-resolution scenario and ResNet-152 [17] for high-resolution scenario, both trained on the FaceScrub. For each attack method, the number of images reconstructed per class is set to 5 due to time and computation cost. More detailed experimental settings are listed in Supp.B.
Notably, we limit the evaluation exhibited in the Experiment section to merely three metrics, including Accuracy, Feature Distance, and FID [19], while Sample Diversity is presented in the Supp.C.1 due to the paper length limitation. To measure the deviation across the target labels, we introduce error bars for Accuracy and Feature Distance. As FID metric is computed based on all the recovered images, it is not essential to calculate error bars. Moreover, VMI [49] and RLBMI [16] will be further evaluated in Supp. C.6 owing to their excessive need of time (See Supp. C.5). We also conduct a user study in Supp. D.4.
@@ 4.2 Evaluation on Different Attack Methods
In this part, we prepare a unified experimental setting for different MI attack methods to ensure fair comparison. The resolution of private and public datasets is set to , indicating relatively easier scenarios. Comparisons of white-box and black-box MI attacks are presented in Table 3.
Remarkably, the PLGMI [60] and LOKT [34] achieve state-of-the-art comprehensive performance in white-box attacks and black-box attacks respectively, showing significant superiority in Accuracy and Feature Distance metrics. However, the lowest FID scores occur in the PPA [43] and C2FMI [59] respectively instead of the above methods. We infer that this is because PPA and C2FMI employ more powerful generators (e.g. StyleGAN2 [24]) as the GAN prior compared to PLGMI and LOKT, leading to more real image generation. Visualization results displayed in Fig 2 further validate the inference.
Notably, the visualization of Mirror(black) in Fig.2 and Fig.6 (in Supp.C.7) is anomalous compared to other methods. We perform further investigation on this special phenomenon in Supp.D.3. The experimental results demonstrate that this phenomenon is highly likely attributable to the genetic algorithm compromising the integrity of StyleGAN’s latent vectors, consequently inducing a deviation in image generation from the normal distribution.
@@ 4.3 Evaluation on Higher Resolution
The results imply the significant impact of GAN priors when attacking private images with higher resolution. All the methods that employ stronger GAN priors maintain low FID scores, including Mirror, C2FMI, and PPA, while other methods suffer from significant degradation in image reality. Despite the primary metric for evaluating MI attacks is Accuracy, the reality of reconstructed images should be ensured within a reasonable range for better image quality. Thus, it is imperative to explore more complex GAN priors with enhanced performance in future research, extending the MI field to more challenging and practical applications. More corresponding visualization results are displayed in Supp.C.7.
@@
4.4 Evaluation on Varied Model Predictive Power
The predictive power of the target model is a crucial factor in determining the effectiveness of MI attacks. Previous work [63] has conducted preliminary experimental validations on simple networks (e.g. LeNet [28]), demonstrating that the performance of the first GAN-based attack GMI is influenced by the predictive power of the target model. Therefore, we evaluate the state-of-the-art MI attacks on target models with varied predictive power to further validate the consistency of this characteristic in the recent attacks. The resolution of datasets is set to and the evaluation results are shown in Fig. 3.
The comparison in Fig. 3 reveals most MI attacks maintain the trend that higher predictive power contributes to better attack performance, which is consistent with the aforementioned characteristic. Specifically, the earlier attack methods (e.g., GMI and KEDMI) presents more fluctuation on the trend across different predictive power, while the recent attack methods (e.g., PLGMI and PPA) show in more stable trend when predictive power increases. This indicates the predictive power of target models plays a crucial role in measuring the performance of MI attacks. Thus one can expect lower privacy leakage in the robust model by balancing the accuracy-privacy trade-off.
@@ 4.5 Evaluation on Defense Strategies
This analysis focuses on the robustness of MI attacks when applied to models with defenses. Notably, we select the first 100 classes subset from FaceScrub as the target dataset due to the time cost. The assessment results are listed in Fig.4 and Fig.5. The configuration of defenses is set following the official parameters, as detailed in Table 9 in the Supplementary.
From Fig.4, the TL [20] achieves the state-of-the-art decrease in Accuracy. However, advanced attack methods have overcome current defense strategies to some extent, such as PLG and PPA. Additionally, some older defense strategies (e.g. MID [52]) are no longer effective against the latest attacks. As for Table 5, the SSD [65] is especially effective when defending against black-box and label-only attacks. This indicates that employing defense in the inference stage is more effective than in the training process.
From Fig.4, we observe that LS [44] exhibits unexpectedly poor performance in metric while it was recently published in the top-tier conference. The potential reason might be the utilized target classifier with relatively high test accuracy, as validated in the Supp.D.1. Furthermore, we conduct further experiments with PPA as the attack method against ResNet-152 trained on high resolution (224×224) scenario, proving that this phenomenon can be extended into other defenses. The results are listed in the Table 5, demonstrating that all the train-time defenses become invalid even with the recommended parameters. More in-depth evaluation on defenses is exhibited in Supp.D.2.
Combining this phenomenon with the above experiment on predictive power, our empirical analysis indicates that the leaked information is strongly correlated to the model prediction accuracy, and current defenses cannot effectively reduce the privacy information without sacrificing of model performance. Our findings emphasize that more reliable and stable defense strategies should be studied. A potential solution is designinga more robust model architecture instead of merely adding a regularizer in the training process.
4.6 Evaluation on Adversarial Robustness
Most defense methods are designed to focus solely on countering MI attacks, while neglecting protection against other types of attacks. However, since both training-time and inference-time defenses can influence the model’s outputs, it becomes imperative to evaluate the robustness of defense-protected models when confronted with adversarial attacks. This necessity highlights the importance of comprehensive security assessments to ensure that enhanced privacy protections do not inadvertently compromise the model’s resilience against diverse attacks.
We select the typical adversarial attacks FGSM [14], PGD [31], BIM [26], and OnePixel [45] as representative to evaluate different MI defenses. The target models are set to IR- trained under multiple defenses on the high-resolution FaceScrub dataset. The results are shown in Table 6. In general, the inference-time defenses demonstrate extraordinary enhancement in adversarial robustness, while train-time defenses fail to improve adversarial robustness. Especially, most train-time defenses reduce adversarial robustness under the untargeted scenario of OnePixel.
Our findings reveal that train-time defenses fail to adequately address multi-faceted robustness requirements. A promising solution is integrating adversarial training into the train-time defense process to enhance robustness.
@@ 4.6 Evaluation on Adversarial Robustness
| NO Defense | - | 0.985 | 0.972 | 0.990 | 307.714 | 0.588 | 50.259 |
| MID | 0.967 | 1.000 | 1.000 | 273.687 | 0.517 | 49.239 | |
| 0.956 | 0.990 | 1.000 | 276.889 | 0.526 | 51.227 | ||
| BiDO | 0.980 | 0.986 | 0.996 | 306.827 | 0.554 | 50.943 | |
| 0.974 | 0.968 | 0.996 | 320.191 | 0.594 | 50.132 | ||
| LS | 0.967 | 0.971 | 0.996 | 281.956 | 0.524 | 38.924 | |
| 0.964 | 0.949 | 0.995 | 302.442 | 0.563 | 39.185 | ||
| TL | 0.976 | 0.994 | 1.000 | 282.394 | 0.513 | 53.023 | |
| 0.975 | 0.982 | 0.996 | 306.528 | 0.561 | 51.489 | ||
| TTS | 0.959 | 0.861 | 0.967 | 143.637 | 0.706 | 41.243 | |
| 0.968 | 0.857 | 0.968 | 147.672 | 0.707 | 42.002 |
| Method | Test Acc | Untargeted Attacks | Targeted Attacks | ||||||
| FGSM | PGD | BIM | OnePixel | FGSM | PGD | BIM | OnePixel | ||
| No defense | 0.982 | 0.997 | 0.995 | 1.000 | 0.665 | 0.087 | 0.170 | 0.161 | 0.000 |
| MID | 0.982 | 1.000 | 0.997 | 0.997 | 0.695 | 0.087 | 0.139 | 0.153 | 0.002 |
| BiDO | 0.940 | 1.000 | 1.000 | 1.000 | 0.826 | 0.085 | 0.114 | 0.112 | 0.007 |
| LS | 0.980 | 0.997 | 0.997 | 0.997 | 0.665 | 0.092 | 0.158 | 0.148 | 0.007 |
| TL | 0.953 | 1.000 | 1.000 | 1.000 | 0.873 | 0.117 | 0.175 | 0.170 | 0.007 |
| TTS | 0.973 | 0.826 | 0.985 | 0.992 | 0.100 | 0.034 | 0.497 | 0.502 | 0.000 |
| Purifier | 0.946 | 0.939 | 0.897 | 0.934 | 0.075 | 0.002 | 0.060 | 0.068 | 0.000 |
| SSD | 0.950 | 0.902 | 0.995 | 0.987 | 0.209 | 0.004 | 0.087 | 0.061 | 0.000 |
@@ 5 Conclusion
In this paper, we develop MIBench, a comprehensive, unified, and reliable benchmark, and provide an extensible and reproducible toolbox for researchers. To the best of our knowledge, this is the first benchmark and first open-source toolbox in the MI field. Our benchmark encompasses 19 of the state-of-the-art MI attack methods and defense strategies, and more algorithms will be continually updated. Based on the implemented toolbox, we establish a consistent experimental environment and conducted extensive evaluation to facilitate fair comparison between different methods. With in-depth analyses, we have benchmarked all the typical MI attacks and defenses under the unified setups and identified new insights for future study.
Societal Impact and Ethical Considerations. A potential negative impact could be malicious users leveraging the implemented attack methods to reconstruct private data from the public system. To alleviate potential dilemma, a cautious approach is to establish access permissions.
@@ A.1 Details of Data Processing
This section introduces three types of dataset processing implemented by our toolbox, including data pre-processing, dataset splitting, and dataset synthesis.
@@
FaceScrub [33]. In the low-resolution version, the original data is center cropped at and then scaled, and the high-resolution is scaled directly.
AFHQ-Dog [4]. The images will be resized to the corresponding resolution.
Stanford Dogs [6]. We crop the images following the official annotations and resize them to the corresponding resolution.
@@
This step is used to model training, including classifiers and conditional generators. For labeled datasets like CelebA and FaceScrub, we use 80% data for training, 10% for validation, and 10% for testing. The validation set is used to select the best trained models, hyperparameters for training, and hyperparameters for defenses. For unlabeled datasets such as FFHQ and MetFaces, we provide a script for pseudo-labeling all images by computing the scores. The strategy of PLGMI [60] for the score calculation is contained in the examples of our toolbox.
@@
Some methods use synthetic datasets for model training. We provide a script to synthesis datasets by pre-trained GANs according to LOKT [34].
@@
By modifying the architecture of the generators and discriminators following PPA [43], we implemented and versions for all the custom attack models defined by the official code of implemented algorithms. For those algorithms that use StyleGAN2 [24], we provide a wrapper for loading the official model and adapting it to our attack process.
@@ A.3 Implementation of Classifier Training & Defense Methods
MID [52]. It applies a Gaussian perturbation to the features in front of the classification head to reduce the mutual information of the model inputs and outputs. According to the official code, the method can also be called VIB, and we adopt this name in our toolbox. The hyperparameter controls the factor of the Gaussian perturbation term.
@@
BiDO [37]. It adds the regular term loss (COCO or HSIC) function so that the intermediate features decrease the mutual information with the input and increase the mutual information with the output. It has two hyperparameters, and . The former is the factor of the regular term loss between inputs and features, and the latter means that between features and model outputs.
@@
LS [44]. It adds a label smoothing term with a negative smoothing factor into the cross-entropy loss.
@@
TL [20]. This method uses a pre-trained model for fine-tuning. The parameters of the previous layers are frozen and those of the later layers will be fine-tuned. In our implementation, the hyperparameter is defined as a ratio of the number of frozen layers.
6.
TTS [25]. TTS first train classifier without defense. After that, the residual shortcuts in the last stages will be removed. Then, the classifier will be continuously trained. The hyperparameter means the keep ratio of a residual shortcut, and denotes the number of residual shortcuts to be removed.
7.
Purifier [57]. Despite the lack of open code and details about and NN, we reproduce it according to their paper. The denotes the weight of keeping the first label when training the CVAE. If the L2 distance between the output and the nearest members, we swap the first and second labels. The validation set is used as the reference set it needs.
8.
SSD [65]. It modifies outputs by solving convex optimization problems. is used for probability calibration and is the distortion bound. The validation set is used to estimate .